OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 2x − 3y = 3 and passes through the point (−8, 2).

OpenStudy (anonymous):

the product of slopes of perpendicular lines is -1 So you know the slope of given line, find the slope of its perpendicular line and use y-y1 = m(x-x1)

OpenStudy (anonymous):

Hint: m1*m2 = -1

OpenStudy (anonymous):

can you work this one out I work better when I have examples and I have more questions I have to do. Please

OpenStudy (anonymous):

slope of give line = -a/b i.e m=2/3 therefore slope of perpndicular on the given line is = -3/2 passing through (-8,2). Now apply slope-point form.